# From Forecasting to Planning: Policy World Model for Collaborative State-Action Prediction

### v1

@@ 4.2 Overall Comparison
On the nuScenes dataset, the relatively simple driving scenarios tend to induce an over‑reliance on the vehicle’s ego status [68, 74]. Therefore, in Table 1, we compare planning performance both without and with access to ego status across several popular methods. PWM achieves the lowest average collision rates (%) of 0.07 and 0.04 across two settings, respectively, surpassing previous state-of-the-art models including Drive-OccWorld [10] and DiffusionDrive [17], with the former adopting a decoupled world modeling paradigm. On the more challenging NAVSIM dataset, as shown in Table 2, PWM delivers obvious superiority. Although we rely on only a single front camera view, our framework significantly outperforms all prior camera-based models, including world modeling methods such as DrivingGPT and LAW. It achieves a PDMS score of 88.1, comparable to the state-of-the-art method DiffusionDrive, which uses both camera and LiDAR inputs. Meanwhile, our model achieves a higher time-to-collision (TTC) score of 95.4 and a no-at-fault collision (NC) score of 98.6.
@@ 4.3 Ablation Study
Impact of Action-free Video World Knowledge. We investigate the influence of learning from large-scale videos on planning performance in Table 3(a) and  3(b). The first rows of the two tables show the results without pre-training, where models are fine-tuned on the downstream task using the base model’s weights [55]. All subsequent rows report the results after pre-training. It shows that without pre-training, the model struggles to capture and predict dynamic scene changes and yields the worst prediction and planning metrics. After pre-training on action-free video generation, the model significantly improves its ability to predict future frames and achieves substantial gains in the planning task. In Table 5, we further compare the model’s performance under different scales of video data on nuScenes. Without pre-training, adding the future-frame prediction task actually harms the model’s planning capability, resulting in much worse performance than directly outputting trajectory waypoints. As the model is exposed to more data, from 0% to 50% and then to 100%, it progressively learns spatiotemporal modeling from driving videos, and the performance gap gradually narrows.

### v2

@@ 4.2 Overall Comparison
On the nuScenes dataset, the relatively simple driving scenarios tend to induce an over‑reliance on the vehicle’s ego status [ego_status, ego_status2]. Therefore, in Table 1, we compare planning performance both without and with access to ego status across several popular methods. PWM achieves the lowest average collision rates (%) of 0.07 and 0.04 across two settings, respectively, surpassing previous state-of-the-art models including Drive-OccWorld [drivinginoccworld] and DiffusionDrive [diffusiondrive], with the former adopting a decoupled world modeling paradigm. On the more challenging NAVSIM dataset, as shown in Table 2, PWM delivers obvious superiority. Although we rely on only a single front camera view, our framework significantly outperforms all prior camera-based models, including world modeling methods such as DrivingGPT and LAW. It achieves a PDMS score of 88.1, comparable to the state-of-the-art method DiffusionDrive, which uses both camera and LiDAR inputs. Meanwhile, our model achieves a higher time-to-collision (TTC) score of 95.4 and a no-at-fault collision (NC) score of 98.6.
@@ 4.3 Ablation Study
Impact of Action-free Video World Knowledge. We investigate the influence of learning from large-scale videos on planning performance in Table 3(a) and  3(b). The first rows of the two tables show the results without pre-training, where models are fine-tuned on the downstream task using the base model’s weights [showo]. All subsequent rows report the results after pre-training. It shows that without pre-training, the model struggles to capture and predict dynamic scene changes and yields the worst prediction and planning metrics. After pre-training on action-free video generation, the model significantly improves its ability to predict future frames and achieves substantial gains in the planning task. In Table 5, we further compare the model’s performance under different scales of video data on nuScenes. Without pre-training, adding the future-frame prediction task actually harms the model’s planning capability, resulting in much worse performance than directly outputting trajectory waypoints. As the model is exposed to more data, from 0% to 50% and then to 100%, it progressively learns spatiotemporal modeling from driving videos, and the performance gap gradually narrows.